OpenStudy (anonymous):

solve and check the inequality if possible. 6|11m-11| greater than or equal to 66

OpenStudy (anonymous):

Its doesn't have a equal sign to the inequality

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[6|11m-11|\ge66\]

OpenStudy (anonymous):

6 | 11m - 11| >= 66 --- divide both sides by 6 | 11m - 11 | >= 11 (11m - 11) >= 11 -- add 11 to both sides 11m >= 11 + 11 11m >= 22 -- divide by 11 m >= 2 (11m - 11) <= -11 -- add 11 to both sides 11m <= 0 m <= 0 solution is : m >= 2 and m <= 0

OpenStudy (anonymous):

This should help ... http://www.purplemath.com/modules/absineq.htm

OpenStudy (anonymous):

thank u son much!!!!

OpenStudy (anonymous):

soo*

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

but how do i check my answer?

OpenStudy (anonymous):

| 11m - 11 | >= 11....we have established that m >= 2, so sub in 3 for m | 11(3) - 11 | >= 11 | 33 - 11 | >= 11 | 22 | >= 11 22 >= 11 (correct) | 11m - 11 | >= 11...we have established that m <=0, so sub in 0 for m | 11(0) - 11 | >= 11 | 0 - 11 | >= 11 | -11 | >= 11 -- remember that all absolute values are positive 11 >= 11 (correct)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

Omg , im so sorry i was suppose to help but ii was doing some else . But Glad u found the answer